OpenStudy (anonymous):

Solve using the square root method. Round the solution to the nearest tenth. x^2 - 20 = 0 A. {-10.0, 10.0} B. {-5.5, 5.5} C. {-3.5, 3.5} D. {-4.5, 4.5}

OpenStudy (anonymous):

x^2=20 \[\Large x=\pm \sqrt{20}\]

OpenStudy (anonymous):

So . D.

OpenStudy (anonymous):

x= 4.472...

Directrix (directrix):

x^2 - 20 = 0 x^2 = 20 x^2 = (2*2*5) Taking square roots of both sides x = ± 2√5 = ± 4.47 approx = ± 4.5 approx --> Option D